OpenStudy (anonymous):

A regular octagon has perimeter of 87.2 in and the area is 573.67 in2. A second octagon has an area equal to 2294.66 in2. Find the perimeter of the second octagon. Round to the nearest hundredth

OpenStudy (anonymous):

All regular \(n\)-gons are similar to other regular \(n\)-gons. This means you can set up a ratio due to the similarity: \[\frac{\text{area}}{\text{perimeter}}=\frac{573.67}{87.2}=\frac{2294.66}{x}\] Solve for \(x\), which is the perimeter of the second octagon.

OpenStudy (anonymous):

Thank you